OpenStudy (anonymous):

The density of matter in interstellar space amounts to 1 atom (H not H2) for every cm^3. (a) What is the pressure of interstellar space assuming a temperature of 4 K?

OpenStudy (aaronq):

PV=nRT; n=m/M -> PV=(m/M)*RT -> P=[m/(V*M)]*RT; d=m/V P=dRT/M

OpenStudy (frostbite):

@aaronq I'm not sure I follow? You choose to assume that the hydrogen atom is an ideal gas?

OpenStudy (aaronq):

yeah, mainly because the pressure is really low, although the temp is not ideal.

OpenStudy (frostbite):

Right. Make since.

OpenStudy (aaronq):

(Y) <- thumbs up

OpenStudy (frostbite):

░░░░░░░░░░░████ ░░░░░░░░░░░█░░█ ░░░░░░░░░░█░░░█ ░░░░░░░░░█░░░░█ ███████▄▄█░░░░░██████▄ ▓▓▓▓▓▓█░░░░░░░░░░░░░░█ ▓▓▓▓▓▓█░░░░░░░░░░░░░░█ ▓▓▓▓▓▓█░░░░░░░░░░░░█ ▓▓▓▓▓▓█░░░░░░░░░░░░░█ ▓▓▓▓▓▓█░░░░░░░░░░░░░░█ ▓▓▓▓▓▓█████░░░░░░████ ███████░░░███████ FROSTBITE APPROVED

OpenStudy (aaronq):

hahaha thats awesome